# Generalization vs Specialization under Concept Shift

### v1

@@ 3.2 Anisotropic features
To study the effects of anisotropy, we consider a two-scale model in which the spectral density of the covariance matrix is a mixture of two point masses at $s_{-}$ and $s_{+}$ with weights $\rho_{-}$ and $\rho_{+}\!=\!1\!-\!\rho_{-}$, respectively. These two modes define subspaces into which we decompose the coefficients, $\beta\!=\!\beta_{-}+\beta_{+}$ with $\Sigma\beta_{\pm}\!=\!s_{\pm}\beta_{\pm}$, and similarly for $\tilde{\beta}$. We write down the bias contribution to prediction risk
@@
where $g_{\pm}(z)\!\equiv\!(m(z)s_{\pm}-z)^{-1}$ and $\pi_{\pm}\!\equiv\!\beta_{\pm}^{\mkern-1.5mu\mathsf{T}}\Sigma\beta_{\pm}/\beta^{\mkern-1.5mu\mathsf{T}}\Sigma\beta$ denotes the signal fraction at each scale. Similarly to the isotropic case, we quantify concept shift using coefficient alignments $\cos\theta_{\pm}\!\equiv\!\beta_{\pm}\!\cdot\!\tilde{\beta}_{\pm}/\|\beta_{\pm}\|\|\tilde{\beta}_{\pm}\|$ and scaling factors $\kappa_{\pm}\!\equiv\!\|\tilde{\beta}_{\pm}\|/\|\beta_{\pm}\|$.
@@
Figure 2 illustrates prediction risk under concept shift for two-scale covariates. We tune the ridge regularization strength such that the in-distribution risk is minimized. We consider the cases where concept shift affects only low-variance features via either $\kappa_{-}$ (Fig 2a) or $\theta_{-}$ (Fig 2b), or only high-variance ones via either $\kappa_{+}$ (Fig 2c) or $\theta_{+}$ (Fig 2d). In all cases, we see that test performance develops nonmonotonic data dependence as test data deviates from in-distribution settings; however, its behavior is qualitatively distinct for low and high-variance concept shift. The region in which more data can hurt occurs at intermediate to high $N/P$ for low-variance concept shift (Fig 2a & b) whereas this region corresponds to low $N/P$ for high-variance concept shift (Fig 2c & d).
@@
This seemingly surprising behavior arises from the fact that it takes more data to learn low-variance features and their effects. At low $N$, high-variance features dominate prediction risk; more data hurts when these features do not adequately predict the response at test time, Fig 2c & d. On the other hand, low-variance features affect test performance only when the training sample size is large enough. Sufficiently strong concept shift on these features thus results in detrimental effects of more data at larger $N$ (compared to high-variance concept shift), Fig 2a & b.
Figure 2: Concept shift generalization exhibits sample size nonmonotonicity for anisotropic features. We illustrate prediction risk for the two-scale model, Eq (11), with aspect ratio $s_{-}/s_{+}\!=\!0.1$, spectral weights $\rho_{+}\!=\!\rho_{-}\!=\!\nicefrac{{1}}{{2}}$ and signal fraction $\pi_{+}\!=\!\pi_{-}\!=\!\nicefrac{{1}}{{2}}$. Concept shift is parametrized by coefficient alignments $\cos\theta_{\pm}\!=\!\beta_{\pm}\!\cdot\!\tilde{\beta}_{\pm}/\|\beta_{\pm}\|\|\tilde{\beta}_{\pm}\|$ and scaling factors $\kappa_{\pm}\!=\!\|\tilde{\beta}_{\pm}\|/\|\beta_{\pm}\|$ where the subscripts indicate the variance $s_{\pm}$ of the affected features. We consider the settings in which concept shift affects either low or high-variance features and either alignment or scale; that is, we vary only one out of four parameters, $\theta_{\pm}$ and $\kappa_{\pm}$, at a time (see legend). a Shrinking coefficients of low-variance features. b Rotating coefficients of low-variance features. c-d Same as (a) & (b) but for high-variance concept shift. Here $\textsc{snr}\!=\!1$ and regularization is in-distribution optimal.
@@ 4 Classification experiments
Although our theory is based on ridge regression, it likely captures some general phenomena of learning under concept shift. Here we consider standard multinomial logistic regression for MNIST [27] and FashionMNIST [28] (see, Appendix C for more details). We modify the test datasets by computing PCA on the images and designating the resulting features as either ‘robust’ or ‘nonrobust’. We shuffle nonrobust features across data points to decorrelate them from the labels while preserving marginal statistics, whereas robust features are unchanged. In Fig 3, we let lower-variance features be nonrobust and depict the test accuracy as a function of the variance threshold that separates robust from nonrobust features. We see that when concept shift is strong and few features are robust, more data hurts test accuracy, in qualitative agreement with our results for ridge regression (cf. Fig 1).
@@ 5 Conclusion and outlook
We introduce a ridge regression model for concept shift. Our model is exactly solvable in the high dimensional limit. We show that concept shift can change the qualitative behavior of generalization performance; for sufficiently strong concept shift, ridge regression fails to generalize even with infinite data. In addition, we demonstrate that input anisotropy can lead to nonmonotonic data dependence of prediction risk. Our results are for optimally tuned ridge regression and thus differ from risk nonmonotonicity due to double descent which is absent under optimal regularization [24].
Taken together, our work provides a fresh perspective on a lesser-studied mode of distribution shift. Our model offers a relatively simple setting for building intuitions and testing hypotheses about concept shift generalization. Our classification experiments suggest that the insights from our theory may apply in more general settings. We hope our work will help initiate systematic investigations of the rich learning phenomena induced by concept shift.
@@ Appendix A Prediction risk in the thermodynamic limit
In the thermodynamic limit—$N,P\to\infty$ and $P/N\to\gamma\in(0,\infty)$—the above traces become deterministic (see Appendix B), and the bias and variance converge to (see also Ref [26])
@@ Appendix B Spectral convergence for random matrix traces
Appendix C Classification experiment details
In Fig 3, we train a standard multinomial logistic regression on MNIST and FashionMNIST datasets, using Adam optimizer with a minibatch size of $500$ and a learning rate of $0.001$ for 2000 epochs. We choose training data points at random (without replacement) to vary training sample size $N$; all of the training data is used when N = 60,000. The variation due to stochastic training and training data construction is smaller than the marker size in Fig 3.

### v2

@@ 3.2 Anisotropic features
To study the effects of anisotropy, we consider a two-scale model in which the spectral density of the covariance matrix is a mixture of two point masses at $s_{-}$ and $s_{+}$ with weights $\rho_{-}$ and $\rho_{+}\!=\!1\!-\!\rho_{-}$, respectively. Without loss of generality, we let $s_{+}\!\geq\!s_{-}$ throughout this section. These two modes define subspaces into which we decompose the coefficients, $\beta\!=\!\beta_{-}+\beta_{+}$ with $\Sigma\beta_{\pm}\!=\!s_{\pm}\beta_{\pm}$, and similarly for $\tilde{\beta}$. We write down the bias contribution to prediction risk
@@
where $g_{\pm}(z)\!\equiv\!(m(z)s_{\pm}-z)^{-1}$ and $\pi_{\pm}\!\equiv\!\beta_{\pm}^{\mkern-1.5mu\mathsf{T}}\Sigma\beta_{\pm}/\beta^{\mkern-1.5mu\mathsf{T}}\Sigma\beta$ denotes the signal fraction at each scale. Similarly to the isotropic case, we quantify concept shift using coefficient alignments $\cos\theta_{\pm}\!\equiv\!\beta_{\pm}\!\cdot\!\tilde{\beta}_{\pm}/\|\beta_{\pm}\|\|\tilde{\beta}_{\pm}\|$ and scaling factors $\kappa_{\pm}\!\equiv\!\|\tilde{\beta}_{\pm}\|/\|\beta_{\pm}\|$, both of which now depend also on variance.
@@
Figure 4 illustrates prediction risk under concept shift for two-scale covariates. We numerically tune the ridge regularization strength $\lambda$ such that the in-distribution risk is minimized and the divergences associated with multiple descent phenomena are completely suppressed. We consider the cases where concept shift affects only low-variance features via either $\kappa_{-}$ or $\cos\theta_{-}$ (Fig 4a and b), or only high-variance ones via either $\kappa_{+}$ or $\cos\theta_{+}$ (Fig 4c and d). In all cases, we see that test performance develops nonmonotonic data dependence as test data deviates from in-distribution settings; however, its behavior is qualitatively distinct for low and high-variance concept shift. To isolate the effects of covariate statistics, we set the signal fraction to $\pi_{+}\!=\!\pi_{-}\!=\!\frac{1}{2}$ so that low and high-variance features carry the same signal strength.
In Fig 4a and b, we depict the effects of concept shift on low-variance features. Panel a corresponds to the case where the low-variance coefficients $\beta_{-}$ shrink at test time, $\kappa_{-}\!\leq\!1$, thus suppressing the signal associated with low-variance features. Similarly to the isotropic case (Fig 2a), prediction risk decreases with $N$ for weak concept shift. However, as $\|\beta_{-}\|$ becomes smaller, we see that generalization performance exhibits nonmonotonic behavior; too much training data can worsen generalization. Panel b turns to the case in which $\beta_{-}$ rotates at test time, $\cos\theta_{-}\!\leq\!1$ (cf. Fig 2b). We observe a similar crossover in the data dependence of prediction risk as concept shift becomes stronger. While generalization improves with more data in the low-data limit, this improvement continues monotonically only for adequately weak concept shifts.
In Fig 4c and d, concept shift affects only high-variance features via shrinking coefficients (panel a) and rotated coefficients (panel b). We see again that strong concept shift leads to nonmonotonic data dependence of prediction risk. However, strong concept shift on high-variance features results in risk maximum at intermediate training data size $N$. This behavior contrasts sharply with low-variance concept shift, depicted in Fig 4a and b.
@@
Indeed we could have anticipated the intriguing differences between concept shift affecting low and high-variance features. The data dependence in Fig 4 results from the fact that it takes more data to learn low-variance features and their effects. At low $N$, high-variance features dominate prediction risk; more data hurts when these features do not adequately predict the response at test time, Fig 4c and d. On the other hand, low-variance features affect test performance only when the training sample size is large enough to influence the model. Sufficiently strong concept shift on these features thus results in detrimental effects of more data at larger $N$ (compared to high-variance concept shift), Fig 4a and b. We emphasize that the nonmonotonic data dependence of prediction risk due to concept shift is unrelated to double descent phenomena (which describe risk nonmonotonicity in suboptimally tuned models).
4 Transformer experiments
To test the applicability of our theoretical framework to realistic scenarios, we train transformers to perform in-context learning (ICL) of (noisy) linear functions, which were argued to perform optimal ridge regression [31, 32, 33]. As many-shot prompting, enabled by recently expanded context windows, has shown promise in improving performance [34], it is timely to investigate how distribution shift affects whether longer context is beneficial.
In our experiment, a transformer takes as its input a series of points $(x_{i},y_{i})$ on an unknown function $y_{i}\!=\!f(x_{i})$ for $i\!=\!1,2,\dots,n-1$, terminating with a ‘query’ $x_{n}$ whose function value $y_{n}$ is the prediction target. The model is trained to minimize the mean square loss $\mathcal{L}\!=\!\mathbb{E}[(y_{n}-\hat{y}_{n})^{2}]$. This setup has proved valuable in developing intuitions about ICL [31, 32, 33, 35, 36, 37, 38]. Here, we focus on noisy linear functions—$f_{\beta}(x)\!=\!\beta^{\mkern-1.5mu\mathsf{T}}x+\xi$ with $\xi$ denoting the noise term—and investigate the generalization properties of the learned ICL solution, implemented by a transformer, under concept shift at test time.
We consider linear regression in $P\!=\!32$ dimensions. The training tasks, parametrized by $\beta$, are drawn iid from a standard normal distribution and kept fixed during training. We generate a total of $2^{20}$ training tasks, which are sufficient to elicit general-purpose ICL [33, 37]. An input sequence is drawn from $y\!=\!\beta^{\mkern-1.5mu\mathsf{T}}x+\xi$ with $(x,\xi)\!\sim\!\mathcal{N}(0,I_{P})\times\mathcal{N}(0,\sigma^{2})$. We choose $\sigma^{2}\!=\!0.5$. The input-response pairs are newly generated each time the model takes an input sequence. We adopt the nanoGPT architecture [39] with eight layers, an embedding dimension of 128, learnable position embeddings, and causal masking. The model is trained to minimize the next token mean squared error (MSE) using Adam with a learning rate of 0.0001. At test time, we sample 10,000 new tasks and compute the in-distribution prediction risk simply as the MSE of the transformer on test tasks. To compute risk under concept shift, the transformer is presented with a context, ($x_{1},y_{1},\dots,x_{n-1},y_{n-1},\tilde{x}$), in which $y_{i}\!=\!\beta^{\mkern-1.5mu\mathsf{T}}x_{i}+\xi$. But the query $\tilde{x}$ is related to the final prediction target $\tilde{y}$ via a linear function $\tilde{y}\!=\!\tilde{\beta}^{\mkern-1.5mu\mathsf{T}}\tilde{x}$ with $\tilde{\beta}\!\neq\!\beta$ in general.
Figure 5: Long context can hurt performance of in-context regression in transformer models. We depict prediction risk of transformers (circles), trained to solve linear regression tasks using in-context examples (see 4 for details). We compare in-context regression We consider two concept shift settings. a Coefficient shrinking parametrized by $\kappa$ (see legend; cf. Fig 2a). b Coefficient rotation parametrized by $\theta$ (see legend; cf. Fig 2b). We compare the asymptotic prediction risk of transformers on isotropic data with the predictions from our theory (Section 3) under two concept shift settings. Circles depict MSE loss attained by the transformer whereas lines depict the optimally regularized asymptotic limit. We observe strong agreement between experimental results and thermodynamic limits. Here, snr=128.
Figure 5 depicts ICL prediction risk for linear regression under concept shift as a function of in-context sample size $N$. We parametrize the degree of concept shift using the scaling factor and cosine similarity, described in Eq (10). We consider two specific settings: in panel a, $\kappa\!\leq\!1$ and $\cos\theta\!=\!1$, and in panel b, $\kappa\!=\!1$ and $\cos\theta\!\leq\!1$ (cf. Fig 2a and b). In both cases, we compare ICL prediction risk (symbols) with our thermodynamic-limit theory (lines). We see that they are in good quantitative agreement. In particular, the context-length dependence ICL prediction risk changes from improving to worsening with longer context as concept shift becomes more severe.
To further investigate how transformers handle concept shift with anisotropic features, we conducted experiments comparing models trained on isotropic versus anisotropic data. Figure 6 illustrates how feature anisotropy influences the generalization behavior of transformer-based in-context regression under varying degrees of concept shift. When exposed to two-scale anisotropic data (see §3.2) where concept shift affects only high-variance features, transformers exhibit markedly different behaviors depending on their training data distribution. Transformers pretrained on anisotropic data closely align with theoretical predictions for optimally tuned ridge regression with structured penalties, indicating that they can apply different penalties to features based on their variances. This strategy is equivalent to whitening the features, resulting in an isotropized regression problem with an effective concept shift scale parameter $\kappa^{2}\!=\!(1+\kappa_{+}^{2})/2$. In contrast, transformers trained on isotropic data struggle to adapt optimally to anisotropic test data, exhibiting more pronounced nonmonotonic risk curves than theoretically predicted for isotropic ridge penalties. This difference becomes particularly evident as concept shift strengthens (panel c), suggesting that the priors learned during pretraining significantly influence how transformers adapt to concept shift in features with different scales. These results highlight the importance of data structure in determining how models respond to distribution shifts and demonstrate that transformers can implicitly learn sophisticated regularization strategies when exposed to appropriately structured training data. We note that this agreement between theory and experiments is not trivial, as transformers perform in-context regression on finite data, in finite dimensions, and via a learned, data-driven algorithm (as opposed to optimal ridge regression in the high-dimensional limit).
Figure 6: Feature anisotropy modulates concept shift effects in transformer-based in-context regression. We depict prediction risk on two-scale, anisotropic data (see §3.2) for optimal ridge regression (curves) and transformers (symbols) trained to solve linear regression tasks using in-context examples. The theoretical results are for optimally tuned isotropic and structured ridge penalties (dashed and solid curves, respectively). The transformers’ results are for isotropic and anisotropic pretraining data (circles and plus symbols, respectively). Concept shift affects only high-variance features via the scaling factor $\kappa_{+}^{2}\!=\!1.0,0.6,0.3$ in a, b, and c, respectively ($\kappa_{-}\!=\!\cos\theta_{\pm}\!=\!1$ throughout). Transformers trained on anisotropic data (pluses) closely follow the solid curve, suggesting that they effectively implement a structured regularization that adapts to feature anisotropy—effectively isotropizing the features and averaging the impact of concept shift ($\kappa^{2}\!=\!(\kappa_{+}^{2}+\kappa_{-}^{2})/2\!=\!(1+\kappa_{+}^{2})/2$). Transformers trained on isotropic data (circles) exhibit nonmonotonic risk under concept shift, but with much more pronounced peaks compared to theoretical predictions (dashed). Here $s_{-}/s_{+}\!=\!0.1$, $\rho_{+}\!=\!\rho_{-}\!=\!\nicefrac{{1}}{{2}}$, $\pi_{+}\!=\!\pi_{-}\!=\!\nicefrac{{1}}{{2}}$ and snr=1.
@@ 5 Classification experiments
[…]
We modify the test datasets by performing principal component analysis (PCA) on the images and designating the resulting features as either ‘robust’ or ‘nonrobust’ (cf. Fig 2c; see also §3.1). We shuffle nonrobust features across data points to decorrelate them from the labels while preserving marginal statistics, whereas robust features are unchanged. In Fig 7, we let lower-variance features be nonrobust and parametrize concept shift strength with the variance threshold that separates robust from nonrobust features. We depict the test accuracy as a function of concept shift strength, and observe that when concept shift is strong and few features are robust, more data hurts test accuracy, in qualitative agreement with our results for linear regression problems (cf. Figs 2 and 5). The qualitative agreement between our theoretical results and experiments on transformers, MNIST, and FashionMNIST suggests that our identification of distinct concept shift regimes provides a conceptual framework applicable to more complex settings.
6 Related works
There is a significant body of work on out-of-distribution generalization with covariate shift, i.e., domain generalization (see, e.g., Refs [8, 42]), but significantly less work has been done on concept shift generalization. However, even within covariate shift, an understanding of out-of-distribution generalization remains elusive [11]. Within concept shift, work has tended to focus instead on detection and mitigation strategies rather than generalization (see, e.g., Ref [29]).
Linear regression has proved a particularly useful setting for investigating learning phenomena. Random matrix universality makes it well-suited for theoretical investigations, yielding much-needed insights into high-dimensional learning [19, 21, 23, 22, 18, 24, 25], including covariate shift generalization [43]. Linear models have also contributed substantially to our developing understanding of in-context learning in transformers [31, 32, 33, 35, 36, 37, 38].
The nonmonotonic behavior we identify is distinct from double descent phenomena [26, 27]. Indeed, our theoretical results in Figs 2-4 are for optimal regularization, for which double descent is absent [17, 28, 18], thus demonstrating that concept shift induces risk nonmonotonicity through fundamentally different mechanisms.
We build on analytical techniques developed for high-dimensional ridge regression, but address a phenomenon distinct from prior works. In particular, optimally tuned models can exhibit nonmonotonic risk due to model misspecification [18]—i.e., a mismatch between model class and data-generating process, irrespective of test-time shift. This mechanism differs from distribution shift, in which the data-generating process itself changes at test time. Our framework specifically isolates the effects of concept shift in correctly specified models.
In-context learning of transformers has been studied under covariate shift [36, 44, 45] but not with respect to concept shift. Song et al. [46] studied how transformers generalize to out-of-distribution tasks through symbolic manipulation via induction heads, but the scenario of within-context concept shift remains underexplored. While Agarwal et al. [34] showed that longer context windows typically benefit in-context learning, they also observed performance degradation in tasks such as MATH.
@@ 7 Conclusion and outlook
We introduce a ridge regression model for concept shift. Our model is exactly solvable in the high dimensional limit. We show that concept shift can change the qualitative behavior of generalization performance; for sufficiently strong concept shift, ridge regression fails to generalize even with infinite data (Figs 2 and 3). In addition, we demonstrate that input anisotropy can lead to nonmonotonic data dependence of prediction risk. In particular, too much data can harm generalization (Fig 4a and b) and more data may only improve generalization above a certain threshold (Fig 4c and d). We emphasize that our results are for optimally tuned ridge regression and thus differ from risk nonmonotonicity due to double and multiple descent phenomena which are absent under optimal regularization [28].
Taken together, our work provides a fresh perspective on a lesser-studied mode of distribution shift. Our model offers a relatively simple setting for building intuitions and testing hypotheses about concept shift generalization. Although our theoretical analyses are exact only for ridge regression in the high-dimensional limit, the qualitative conclusions generalize beyond this idealized setting. Indeed, our theoretical prediction agrees quantitatively with experiments on regression in finite dimensions from finite samples, using the in-context learning ability of a transformer model as a learning algorithm. Additionally, our classification experiments suggest that the insights from our theory may apply in more general settings. Finally, our theoretical results for anisotropic features have implications for understanding data dimension reduction techniques—such as principal component regression in which low-variance features are discarded—under concept shift. We hope our work will encourage further systematic investigations of the rich learning phenomena induced by concept shift.
Acknowledgments and Disclosure of Funding
Alex Nguyen is supported by NIH grant RF1MH125318. DJS was partially supported by a Simons Fellowship in the MMLS, a Sloan Fellowship, and the National Science Foundation, through the Center for the Physics of Biological Function (PHY-1734030). VN acknowledges research funds from the University of Sydney.
@@ Appendix A Prediction risk in the thermodynamic limit
In the thermodynamic limit—$N,P\to\infty$ and $P/N\to\gamma\in(0,\infty)$—the above traces become deterministic (see Appendix B), and the bias and variance converge to (see also Patil et al. [30])